# An Efficient Row-Based Sparse Fine-Tuning

### v1

@@ Appendix A Importance Metrics
To prove Property 4.2, we first calculate the expectation and variance of SPSA. For convenience, we denote $[\nabla L(\mathbf{\theta},\mathcal{D})]$ as ${\mathbf{g}}$. Then, the expectation is as follows:
$\mathbb{E}[\hat{{\mathbf{g}}}]$ $\approx\mathbb{E}[\mathbf{z}\mathbf{z}^{\top}\nabla L(\mathbf{\theta},\mathcal{D})]=\mathbb{E}[\mathbf{z}\mathbf{z}^{\top}]\nabla L(\mathbf{\theta},\mathcal{D})=\mathbf{I}_{d}\nabla L(\mathbf{\theta},\mathcal{D}).$
@@
Thus, the expectation and variance of $n$-SPSA estimate are
$\mathbb{E}[\frac{\sum_{j=1}^{n}\hat{{\mathbf{g}}}_{i}^{(j)}}{n}]$ $=\frac{\sum_{j=1}^{n}\mathbb{E}[\hat{{\mathbf{g}}}_{i}^{(j)}]}{n}=\frac{\sum_{j=1}^{n}{\mathbf{g}}_{i}}{n}={\mathbf{g}}_{i}.$
@@
which converges to 00 as $n\rightarrow\infty$.
Appendix B Parameter Dependency
Dependencies of parameters between neurons or channels across different layers exist in NNs. These include basic layer connections, residual connections, tensor concatenations, summations, and more, as shown in Figure 2. The black neurons connected by real lines represent the dependent parameters that are in the same group. Pruning any black neurons results in removing the parameters connected by the real lines. (Liu et al., 2021) introduced a group pruning method for CNN models that treats residual connections as grouped dependencies, evaluating and pruning related channels within the same group simultaneously. Similarly, (Fang et al., 2023) proposed a novel group pruning technique named Torch-Pruning, which considers various types of dependencies and achieves state-of-the-art results. (Ma et al., 2023) further applied this procedure to pruning LLMs. Torch-Pruning can be applied to prune a wide range of neural networks, including image transformers, LLMs, CNNs, and more, making it a popular toolkit for neural network pruning.
@@ Appendix B Parameter Dependency
In this study, we also evaluate the influences of incorporating parameter dependency in our approach. We put the experimental results of whether incorporating parameter dependency in Appendix C.2. In the experiments, parameter dependency becomes the following process for our approach: first, searching for dependencies by tracing the computation graph of gradient; next, evaluating the importance of parameter groups; and finally, fine-tuning the parameters within those important groups collectively. For instance, if ${\mathbf{W}}^{a}_{\cdot j}$ and ${\mathbf{W}}^{b}_{i\cdot}$ are dependent, where ${\mathbf{W}}^{a}_{\cdot j}$ is the $j$-th column in parameter matrix (or the $j$-th input channels/features) of layer $a$ and ${\mathbf{W}}^{b}_{i\cdot}$ is the $i$-th row in parameter matrix (or the $i$-th output channels/features) of layer $b$, then ${\mathbf{W}}^{a}_{\cdot j}$ and ${\mathbf{W}}^{b}_{i\cdot}$ will be fine-tuned simultaneously while the corresponding ${\mathbf{M}}^{a}_{dep}$ for ${\mathbf{W}}^{a}_{\cdot j}$ becomes column selection matrix and ${\mathbf{W}}^{a}_{s}$ becomes ${\mathbf{W}}^{a}_{f,dep}{\mathbf{M}}^{a}_{dep}$. Consequently, fine-tuning $2.5\%$ output channels for layer $b$ will result in fine-tuning additional $2.5\%$ input channels in each dependent layer. Therefore, for the $5\%$ of desired fine-tuning ratio, the fine-tuning ratio with considering dependencies is set to $2.5\%$ [^8] In some complex models, considering dependencies results in slightly more than twice the number of trainable parameters. However, in most cases, the factor is 2. for the approach that includes dependencies.
@@
$f(\hat{{\mathbf{W}}}^{a},{\mathbf{x}})$ $=f({\mathbf{W}}^{a},{\mathbf{x}})+f({\mathbf{M}}^{a}{\mathbf{W}}^{a}_{f},{\mathbf{x}})+f({\mathbf{W}}^{a}_{f,dep}{\mathbf{M}}^{a}_{dep},{\mathbf{x}}).$
Note that in this example, as the dependency is connection between the output feature/channel of $b$ and the input feature/channel of $a$, the dimension $d^{a}_{in}$ is equal to $d^{b}_{out}$ where ${\mathbf{W}}^{a}\in\mathbb{R}^{d^{a}_{out}\times d^{a}_{in}},{\mathbf{W}}^{b}\in\mathbb{R}^{d^{b}_{out}\times d^{b}_{in}}$.
Appendix C Ablation Studies and Related Analysis
@@
C.1 Hyperparameter Settings
We report the results of three approaches over several epochs as table 7 and table 8. Overall, full fine-tuning over higher epochs is more prone to overfitting, while head fine-tuning shows the exact opposite trend. Except for the results on caltech101999The inconsistent trend observed in Caltech101 results is likely due to its significantly smaller sample size., the loss patterns across all models consistently reflect this trend, and most accuracy results further support this conclusion. However, our approach demonstrates a crucial advantage by effectively balancing the tradeoff between performance and computational resources.
@@ C.1 Hyperparameter Settings
We refer from Table 8 for epochs choosing of fine-tuning Llama2 and Llama3. Table 8 shows that 5 or 10 epochs are reasonable for most tasks using our approach. Given that the maximum sequence lengths of Llama are longer than DeBERTaV3, we have opted for only 5 epochs in the main experiments to balance computational resources and performance.
C.2 Considering Dependency
@@ C.3 Memory Measurement
FA-LoRA, $r=64$ 92.8M(1.38%) 47.12GB 29.87GB 17.25GB 113.2M(1.41%) 58.41GB 33.86GB 30.37GB
FA-RoSA, $r=32,d=1.2\%$ 98.3M(1.46%) 68.21GB 29.87GB 38.34GB 124.3M(1.55%) 76.17GB 33.86GB 42.31GB
FA-DoRA, $r=64$ 93.6M(1.39%) 60.48GB 29.87GB 30.61GB 114.3M(1.42%) 74.48GB 33.86GB 40.62GB
FA-SPruFT, $r=128$ 78.6M(1.17%) 45.08GB 29.87GB 15.21GB 92.3M(1.15%) 56.27GB 33.86GB 22.41GB
Table 11: The requirements of computation resources for fine-tuning. ‘mem’ traces $\text{mem}_{\text{TTL}}-\text{mem}_{\text{M}}$. All fine-tuning parameters are stored in full precision. We also examined the training time and observed that DoRA requires 50% to 100% more time than other methods, while LoRA, RoSA, and our approach need similar training time (differing only by a few seconds). However, due to the influence of various factors on training time and the difficulty of ensuring a fair comparison, we chose not to include these results in our report.
@@ C.4 Resource Requirements
Figure 4: The illustration of backpropagation highlights the operations involved. Black operations occur during the forward pass, while orange operations take place during the backward pass. Blue operations highlight the benefits of our approach. Notably, since ${\mathbf{M}}$ is non-trainable, caching ${\mathbf{W}}_{f}{\mathbf{x}}$ during the forward pass is unnecessary, leading to significant memory savings. Additionally, in practice, PyTorch caches $\frac{\partial L}{\partial{\mathbf{h}}_{\text{right}}}$ to efficiently compute $\frac{\partial L}{\partial{\mathbf{B}}}$, although this caching is not strictly required for backpropagation.
C.5 Cache Benefit
@@
In the main context, we have already shown the memory cost of dropout layers in LoRA, in this section, we will discuss some other benefits of our approach. Figure 4 illustrates the computation and cache requirements in backpropagation (Rumelhart et al., 1986). For simplicity, we replace the notation $f(\cdot,\cdot)$ with different ${\mathbf{h}}$. With the same number of trainable parameters, our approach eliminates the need to cache ${\mathbf{h}}={\mathbf{W}}_{f}{\mathbf{x}}$ shown in the figure. While this benefit is negligible under lower rank settings ($r$) or when the number of fine-tuning layers is small, it becomes significant as the model size and rank settings increase. Although the caching requirement for ${\mathbf{h}}$ can be addressed by recomputing ${\mathbf{h}}={\mathbf{A}}{\mathbf{x}}$ during backpropagation, this would result in increased time complexity during training.
Model, ft setting mem #param ARC-c ARC-e BoolQ HS OBQA PIQA rte SIQA WG Avg
Llama2(7B), LoRA, $r=16$ 21.64GB 40.0M(0.59%) 44.71 76.89 77.49 57.94 32.2 78.73 60.65 47.59 68.75 60.55
Llama2(7B), SPruFT, $r=32$ 15.57GB 36.4M(0.54%) 43.34 77.43 77.80 57.06 32.2 78.02 63.18 46.52 69.14 60.52
Llama2(7B), LoRA, $r=32$ 22.21GB 80.0M(1.19%) 44.28 76.89 77.37 57.61 32.0 78.45 64.62 47.90 69.14 60.92
Llama2(7B), SPruFT, $r=64$ 16.20GB 72.9M(1.08%) 43.69 77.30 77.83 57.19 32.2 78.07 63.54 47.03 69.22 60.67
Llama2(7B), LoRA, $r=64$ 23.46GB 159.9M(2.37%) 44.97 77.02 77.43 57.75 32.0 78.45 62.09 47.75 68.75 60.69
Llama2(7B), SPruFT, $r=128$ 17.62GB 145.8M(2.16%) 43.60 77.26 77.77 57.47 32.6 78.07 64.98 46.67 69.30 60.86
Llama2(7B), FA-LoRA, $r=16$ 16.29GB 23.2M(0.34%) 44.54 77.36 77.83 57.39 30.8 77.69 67.15 47.08 68.82 60.96
Llama2(7B), FA-SPruFT, $r=32$ 14.16GB 19.7M(0.29%) 43.69 77.30 77.55 57.14 31.6 78.02 64.62 46.88 69.38 60.69
Llama2(7B), FA-LoRA, $r=32$ 16.56GB 46.4M(0.69%) 44.03 77.48 77.61 57.40 30.4 77.80 65.70 46.98 68.98 60.71
Llama2(7B), FA-SPruFT, $r=64$ 14.48GB 39.3M(0.58%) 43.77 77.26 77.65 57.17 31.8 78.07 65.70 46.78 69.22 60.82
Llama2(7B), FA-LoRA, $r=64$ 17.25GB 92.8M(1.38%) 43.77 77.57 77.74 57.45 31.0 77.86 66.06 47.13 69.06 60.85
Llama2(7B), FA-SPruFT, $r=128$ 15.21GB 78.6M(1.17%) 43.94 77.22 77.83 57.11 32.0 78.18 65.70 46.47 69.38 60.87
Llama3(8B), LoRA, $r=16$ 28.86GB 41.9M(0.52%) 53.50 81.44 82.35 60.61 34.2 79.87 69.31 48.00 73.56 64.76
Llama3(8B), SPruFT, $r=32$ 22.62GB 39.8M(0.50%) 51.19 80.81 81.10 60.21 34.4 79.60 70.40 47.19 73.01 64.21
Llama3(8B), LoRA, $r=32$ 29.37GB 83.9M(1.04%) 53.33 81.86 82.20 60.65 34.0 79.87 68.23 49.03 73.72 64.76
Llama3(8B), SPruFT, $r=64$ 23.23GB 79.7M(0.99%) 51.96 81.27 81.31 60.18 35.2 79.76 70.04 47.80 73.09 64.51
Llama3(8B), LoRA, $r=64$ 30.37GB 167.8M(2.09%) 53.07 81.40 82.32 60.67 34.2 79.98 69.68 48.52 73.56 64.82
Llama3(8B), SPruFT, $r=128$ 24.49GB 159.4M(1.98%) 52.47 81.10 81.28 60.29 34.6 79.76 70.04 47.75 73.24 64.50
Llama3(8B), FA-LoRA, $r=16$ 23.54GB 28.3M(0.35%) 51.45 81.48 82.17 60.17 34.4 79.60 68.95 47.75 73.48 64.38
Llama3(8B), FA-SPruFT, $r=32$ 21.24GB 23.1M(0.29%) 50.60 80.85 81.19 60.22 34.2 79.54 70.76 47.24 73.01 64.18
Llama3(8B), FA-LoRA, $r=32$ 23.89GB 56.6M(0.71%) 52.22 81.61 82.35 60.26 35.0 79.98 69.68 48.41 73.80 64.81
Llama3(8B), FA-SPruFT, $r=64$ 21.62GB 46.1M(0.57%) 51.79 81.06 81.22 60.20 35.2 79.65 71.48 47.24 73.16 64.56
Llama3(8B), FA-LoRA, $r=64$ 24.55GB 113.2M(1.41%) 52.47 81.36 82.23 60.17 35.0 79.76 70.04 48.31 73.56 64.77
Llama3(8B), FA-SPruFT, $r=128$ 22.41GB 92.3M(1.15%) 52.22 81.19 81.35 60.20 34.2 79.71 69.31 47.13 73.01 64.26
@@
Table 12: Similar to Table 3, this is the full table with different rank settings for our SPruFT and LoRA. We also present the results of freezing-attention blocks in this table.
Model, ft setting ARC-c ARC-e BoolQ HS OBQA PIQA rte SIQA WG Avg

### v2

@@ Appendix B Missing Proofs
Proof.
To prove Property 4.2, we first calculate the expectation and variance of SPSA. For convenience, we denote $[\nabla L({\bm{\theta}},\mathcal{D})]$ as ${\mathbf{g}}$. Then, the expectation is as follows:
$\mathbb{E}[\hat{{\mathbf{g}}}]$ $\approx\mathbb{E}[\mathbf{z}\mathbf{z}^{\top}\nabla L({\bm{\theta}},\mathcal{D})]=\mathbb{E}[\mathbf{z}\mathbf{z}^{\top}]\nabla L({\bm{\theta}},\mathcal{D})=\mathbf{I}_{d}\nabla L({\bm{\theta}},\mathcal{D}).$
@@
Thus, the expectation and variance of $n$-SPSA estimate are ($\hat{{\mathbf{g}}}_{i}^{(j)}$ is the $j^{th}$ estimate of $\hat{{\mathbf{g}}}_{i}$ here):
$\mathbb{E}[\frac{\sum_{j=1}^{n}\hat{{\mathbf{g}}}_{i}^{(j)}}{n}]$ $=\frac{\sum_{j=1}^{n}\mathbb{E}[\hat{{\mathbf{g}}}_{i}^{(j)}]}{n}=\frac{\sum_{j=1}^{n}{\mathbf{g}}_{i}}{n}={\mathbf{g}}_{i},$
@@
which converges to 00 as $n\rightarrow\infty$. ∎
Proof of Theorem 4.3
Proof.
$\displaystyle p\left(\hat{\bm{\eta}}_{i_{1}}-\hat{\bm{\eta}}_{i_{2}}>0\right)=\displaystyle p_{Z\sim\mathcal{N}(0,1)}\left(Z>-\frac{{\bm{\eta}}_{i_{1}}-{\bm{\eta}}_{i_{2}}}{\sqrt{\left(\frac{\sum_{j\in\{j^{(i_{1})}\}}\mathbf{\sigma}_{j}^{2}{\bm{\theta}}_{j}^{2}}{2}+\frac{\sum_{j\in\{j^{(i_{2})}\}}\mathbf{\sigma}_{j}^{2}{\bm{\theta}}_{j}^{2}}{2}\right)/2}}\right)$ (16)
$\sum_{j=1}^{d_{i}}\mathbf{\sigma}_{j}^{2}{\bm{\theta}}_{j}^{2}=\sum_{j=1}^{d_{i}}\frac{{\mathbf{g}}_{j}^{2}+\sum_{l=1}^{d}{\mathbf{g}}_{l}^{2}}{nk}{\bm{\theta}}_{j}^{2}\leq\frac{\sum_{j=1}^{d_{i}}{\mathbf{g}}_{j}^{2}+d_{i}\sum_{l=1}^{d}{\mathbf{g}}_{l}^{2}}{nk}u_{\bm{\theta}}^{2}\overset{(*)}{\approx}\frac{d_{i}u_{\bm{\theta}}^{2}\sum_{l=1}^{d}{\mathbf{g}}_{l}^{2}}{nk}$ (17)
$\implies$ $\left(\frac{\sum_{j\in\{j^{(i_{1})}\}}\mathbf{\sigma}_{j}^{2}{\bm{\theta}}_{j}^{2}}{2}+\frac{\sum_{j\in\{j^{(i_{2})}\}}\mathbf{\sigma}_{j}^{2}{\bm{\theta}}_{j}^{2}}{2}\right)/2\leq\frac{(d_{i_{1}}+d_{i_{2}})u_{\bm{\theta}}^{2}\sum_{l=1}^{d}{\mathbf{g}}_{l}^{2}}{4nk}\leq\frac{(d_{i_{1}}+d_{i_{2}})u_{\bm{\theta}}^{2}u_{{\mathbf{g}}}^{2}}{4nk}$ (18)
$\implies$ $\displaystyle p\left(\hat{\bm{\eta}}_{i_{1}}-\hat{\bm{\eta}}_{i_{2}}>0\right)\geq\left(Z>-\frac{2\sqrt{nk}({\bm{\eta}}_{i_{1}}-{\bm{\eta}}_{i_{2}})}{\sqrt{(d_{i_{1}}+d_{i_{2}})u_{\bm{\theta}}^{2}u_{{\mathbf{g}}}^{2}}}\right)\overset{(**)}{\geq}1-\Phi\left(-\frac{2\sqrt{nk}(q_{2}-q_{1})/100}{\alpha\sqrt{(d_{i_{1}}+d_{i_{2}})u_{\bm{\theta}}^{2}u_{\mathbf{g}}^{2}}}\right).$ (19)
(*): $d_{i}$ is the dimensionality of neuron $i$ and $d$ is the number of parameters, so we have $d\gg d_{i}$ and, further, $d_{i}\sum_{l=1}^{d}{\mathbf{g}}_{l}^{2}\gg\sum_{j=1}^{d_{i}}{\mathbf{g}}_{j}^{2}$.
(**): ${\bm{\eta}}_{i}$ is $\alpha$-smooth in $[0,u_{\bm{\eta}}]$, so $p_{{\bm{\eta}}_{i}}\left({\bm{\eta}}_{i_{2}}\leq{\bm{\eta}}_{i}\leq{\bm{\eta}}_{i_{1}}\right)=\frac{q_{2}-q_{1}}{100}\leq\alpha({\bm{\eta}}_{i_{1}}-{\bm{\eta}}_{i_{2}})$, implying that ${\bm{\eta}}_{i_{1}}-{\bm{\eta}}_{i_{2}}\geq\frac{(q_{2}-q_{1})/100}{\alpha}$. ∎
Proof of Theorem 4.4
Proof.
We will argue that for any $i\leq i_{\xi}(1-\frac{\epsilon_{\xi}}{2})$, we have $p(X_{i}=1)\geq 1-\frac{\epsilon_{\xi}}{2}$. This implies that
$\sum_{i=1}^{i_{\xi}}p_{i}\geq i_{\xi}\left(1-\frac{\epsilon_{\xi}}{2}\right)^{2}\geq i_{\xi}(1-\epsilon_{\xi}),$ (20)
as desired. Let us thus focus on $p(X_{i}=1)$ for some $i\leq i_{\xi}(1-\frac{\epsilon_{\xi}}{2})$. Using the lower bound on $p_{i}$ from equation 12, it suffices to prove that
$p(\hat{\bf\eta}_{i}>\hat{\bf\eta}_{i_{\xi}})\geq 1-\frac{\epsilon_{\xi}}{2d_{\bm{\eta}}}.$
Thus using Theorem 4.3, it suffices to prove that
$\Phi\left(-\frac{2c_{d_{\bm{\eta}}}\alpha\sqrt{(d_{i}+d_{i_{\xi}})u_{\bm{\theta}}^{2}u_{\mathbf{g}}^{2}}(\epsilon_{\xi}/2)/100}{\alpha\sqrt{(d_{i}+d_{i_{\xi}})u_{\bm{\theta}}^{2}u_{\mathbf{g}}^{2}}}\right)=\Phi\left(-\frac{c_{d_{\bm{\eta}}}\epsilon_{\xi}}{100}\right)\leq\frac{\epsilon_{\xi}}{2d_{\bm{\eta}}}.$ (21)
Given the fact that $\Phi\left(-\frac{c_{d_{\bm{\eta}}}\epsilon_{\xi}}{100}\right)=\displaystyle p_{Z\sim\mathcal{N}(0,1)}\left(Z\leq-\frac{c_{d_{\bm{\eta}}}\epsilon_{\xi}}{100}\right)$, we have the following statement: the greater the $c_{d_{\bm{\eta}}}$ is, the smaller the $\Phi\left(-\frac{c_{d_{\bm{\eta}}}\epsilon_{\xi}}{100}\right)$ is. This concludes that, there must exist a constant $C_{d_{\bm{\eta}}}$ such that $\Phi\left(-\frac{c_{d_{\bm{\eta}}}\epsilon_{\xi}}{100}\right)$ satisfies Equation 21 $\forall c_{d_{\bm{\eta}}}\geq C_{d_{\bm{\eta}}}$. ∎
Appendix C Parameter Dependency
Dependencies of parameters between neurons or channels across different layers exist in NNs. These include basic layer connections, residual connections, tensor concatenations, summations, and more, as shown in Figure 3. The black neurons connected by real lines represent the dependent parameters that are in the same group. Pruning any black neurons results in removing the parameters connected by the real lines. Liu et al. (2021) introduced a group pruning method for CNN models that treats residual connections as grouped dependencies, evaluating and pruning related channels within the same group simultaneously. Similarly, Fang et al. (2023) proposed a novel group pruning technique named Torch-Pruning, which considers various types of dependencies and achieves state-of-the-art results. Ma et al. (2023) further applied this procedure to pruning LLMs. Torch-Pruning can be applied to prune a wide range of neural networks, including image transformers, LLMs, CNNs, and more, making it a popular toolkit for neural network pruning.
@@ Appendix C Parameter Dependency
In this study, we also evaluate the influences of incorporating parameter dependency in our approach. We put the experimental results of whether incorporating parameter dependency in Appendix D.2. In the experiments, parameter dependency becomes the following process for our approach: first, searching for dependencies by tracing the computation graph of gradient; next, evaluating the importance of parameter groups; and finally, fine-tuning the parameters within those important groups collectively. For instance, if ${\mathbf{W}}^{(a)}_{\cdot j}$ and ${\mathbf{W}}^{(b)}_{i\cdot}$ are dependent, where ${\mathbf{W}}^{(a)}_{\cdot j}$ is the $j$-th column in parameter matrix (or the $j$-th input channels/features) of layer $a$ and ${\mathbf{W}}^{(b)}_{i\cdot}$ is the $i$-th row in parameter matrix (or the $i$-th output channels/features) of layer $b$, then ${\mathbf{W}}^{(a)}_{\cdot j}$ and ${\mathbf{W}}^{(b)}_{i\cdot}$ will be fine-tuned simultaneously while the corresponding ${\mathbf{M}}^{(a)}_{dep}$ for ${\mathbf{W}}^{(a)}_{\cdot j}$ becomes column selection matrix and ${\bm{\delta}}^{(a)}$ becomes $\mathrm{\Delta}{\mathbf{W}}^{(a)}_{dep}{\mathbf{M}}^{(a)}_{dep}$. Consequently, fine-tuning $2.5\%$ output channels for layer $b$ will result in fine-tuning additional $2.5\%$ input channels in each dependent layer. Therefore, for the $5\%$ of desired fine-tuning ratio, the fine-tuning ratio with considering dependencies is set to $2.5\%$ [^10] In some complex models, considering dependencies results in slightly more than twice the number of trainable parameters. However, in most cases, the factor is 2. for the approach that includes dependencies.
@@
$f(\hat{{\mathbf{W}}}^{(a)},{\mathbf{x}})$ $=f({\mathbf{W}}^{(a)},{\mathbf{x}})+f({\mathbf{M}}^{(a)}\mathrm{\Delta}{\mathbf{W}}^{(a)},{\mathbf{x}})+f(\mathrm{\Delta}{\mathbf{W}}^{(a)}_{dep}{\mathbf{M}}^{(a)}_{dep},{\mathbf{x}}).$
Note that in this example, as the dependency is connection between the output feature/channel of $b$ and the input feature/channel of $a$, the dimension $d^{(a)}_{in}$ is equal to $d^{(b)}_{out}$ where ${\mathbf{W}}^{(a)}\in\mathbb{R}^{d^{(a)}_{out}\times d^{(a)}_{in}},{\mathbf{W}}^{(b)}\in\mathbb{R}^{d^{(b)}_{out}\times d^{(b)}_{in}}$.
Appendix D Ablation Studies and Related Analysis
@@
D.1 Hyperparameter Settings
We report the results of three approaches over several epochs as table 9 and table 10. Overall, full fine-tuning over higher epochs is more prone to overfitting, while head fine-tuning shows the exact opposite trend. Except for the results on caltech101111111The inconsistent trend observed in Caltech101 results is likely due to its significantly smaller sample size., the loss patterns across all models consistently reflect this trend, and most accuracy results further support this conclusion. However, our approach demonstrates a crucial advantage by effectively balancing the tradeoff between performance and computational resources.
@@ D.1 Hyperparameter Settings
D.2 Considering Dependency
@@ D.3 Memory Measurement
Table 13: The requirements of computation resources for fine-tuning. ‘mem’ traces $\text{mem}_{\text{TTL}}-\text{mem}_{\text{M}}$. indicates that we freeze the layers for queue, key, and value projection. All fine-tuning parameters are stored in full precision. We also examined the training time and observed that DoRA requires 50% to 100% more time than other methods, while LoRA, RoSA, and our approach need similar training time (differing only by a few seconds). However, due to the influence of various factors on training time and the difficulty of ensuring a fair comparison, we chose not to include these results in our report.
@@
LoRA, $r=64$ 109.6M(1.63%) 48.68GB 29.87GB 18.81GB 130.0M(1.62%) 59.90GB 33.86GB 26.04GB
RoSA, $r=32,d=1.2\%$ 113.2M(1.68%) 69.70GB 29.87GB 39.83GB 139.1M(1.74%) 77.60GB 33.86GB 43.74GB
DoRA, $r=64$ 110.5M(1.64%) 64.05GB 29.87GB 34.18GB 131.2M(1.64%) 77.97GB 33.86GB 44.11GB
VeRA, $r=64$ 0.84M(0.01%) 47.63GB 29.87GB 17.76GB 1.05M(0.01%) 59.64GB 33.86GB 25.78GB
SPruFT, $r=128$ 95.4M(1.42%) 45.45GB 29.87GB 15.58GB 109.1M(1.36%) 56.57GB 33.86GB 22.71GB
@@ D.4 Resource Requirements
Figure 5: The illustration of backpropagation highlights the operations involved. Black operations occur during the forward pass, while orange operations take place during the backward pass. Blue operations highlight the benefits of our approach. Notably, since ${\mathbf{M}}$ is non-trainable, caching $\mathrm{\Delta}{\mathbf{W}}{\mathbf{x}}$ during the forward pass is unnecessary, leading to significant memory savings. Additionally, in practice, PyTorch caches $\frac{\partial L}{\partial{\mathbf{h}}_{\text{right}}}$ to efficiently compute $\frac{\partial L}{\partial{\mathbf{B}}}$, although this caching is not strictly required for backpropagation.
D.5 Cache Benefit
@@
In the main context, we have already shown the memory cost of dropout layers in LoRA, in this section, we will discuss some other benefits of our approach. Figure 5 illustrates the computation and cache requirements in backpropagation (Rumelhart et al., 1986). For simplicity, we replace the notation $f(\cdot,\cdot)$ with different ${\mathbf{h}}$. With the same number of trainable parameters, our approach eliminates the need to cache ${\mathbf{h}}=\mathrm{\Delta}{\mathbf{W}}{\mathbf{x}}$ shown in the figure. While this benefit is negligible under lower rank settings ($r$) or when the number of fine-tuning layers is small, it becomes significant as the model size and rank settings increase. Although the caching requirement for ${\mathbf{h}}$ can be addressed by recomputing ${\mathbf{h}}={\mathbf{A}}{\mathbf{x}}$ during backpropagation, this would result in increased time complexity during training.
@@
Table 14: Similar to Table 2, this is the full table with different rank settings for our SPruFT and LoRA. We also present the results of freezing queue-, key-, and value-projection in this table ()
Model, ft setting mem #param BoolQ PIQA SIQA HS WG ARC-c ARC-e OBQA Avg
Llama2(7B), LoRA, $r=16$ 21.64GB 40.0M(0.59%) 76.4 71.4 69.4 81.2 59.8 67.0 81.4 73.2 72.48
Llama2(7B), SPruFT, $r=32$ 15.57GB 36.4M(0.54%) 78.2 73.6 67.4 79.4 58.8 69.0 82.2 69.4 72.25
Llama2(7B), LoRA, $r=32$ 22.21GB 80.0M(1.19%) 74.8 74.2 71.4 78.4 58.6 67.0 82.2 69.6 72.03
Llama2(7B), SPruFT, $r=64$ 16.20GB 72.9M(1.08%) 78.0 73.8 66.4 83.4 59.6 69.4 79.0 71.6 72.65
Llama2(7B), LoRA, $r=64$ 23.46GB 159.9M(2.37%) 77.0 76.2 67.8 84.2 62.6 70.0 82.0 74.0 74.23
Llama2(7B), SPruFT, $r=128$ 17.62GB 145.8M(2.16%) 80.0 75.2 67.6 85.0 63.4 70.8 82.4 71.8 74.53
Llama2(7B), LoRA, $r=16$ 17.62GB 27.4M(0.41%) 80.4 75.6 68.2 77.4 58.0 66.4 80.2 71.8 72.25
Llama2(7B), SPruFT, $r=32$ 14.29GB 23.9M(0.35%) 82.6 73.0 64.4 80.0 61.4 64.8 80.4 70.2 72.10
Llama2(7B), LoRA, $r=32$ 17.95GB 54.8M(0.81%) 75.6 76.0 68.6 79.2 63.6 67.2 82.0 71.0 72.90
Llama2(7B), SPruFT, $r=64$ 14.67GB 47.7M(0.71%) 78.2 75.8 66.4 81.4 59.2 67.8 78.8 72.0 72.45
Llama2(7B), LoRA, $r=64$ 18.81GB 109.6M(1.63%) 79.6 73.8 67.0 80.4 62.6 69.2 81.0 70.4 73.00
Llama2(7B), SPruFT, $r=128$ 15.58GB 95.4M(1.42%) 81.4 75.0 67.2 80.4 65.0 69.2 79.8 73.8 73.98
Llama3(8B), LoRA, $r=16$ 28.86GB 41.9M(0.52%) 85.2 80.8 68.4 81.8 69.0 79.4 90.0 77.0 78.95
Llama3(8B), SPruFT, $r=32$ 22.62GB 39.8M(0.50%) 81.2 82.2 68.2 85.4 63.4 79.4 87.6 80.8 78.53
Llama3(8B), LoRA, $r=32$ 29.37GB 83.9M(1.04%) 85.2 81.8 68.2 87.8 67.0 76.4 89.2 80.4 79.50
Llama3(8B), SPruFT, $r=64$ 23.23GB 79.7M(0.99%) 83.2 80.6 69.4 86.0 60.6 78.4 83.4 81.4 77.88
Llama3(8B), LoRA, $r=64$ 30.37GB 167.8M(2.09%) 84.2 77.0 63.2 84.2 67.2 76.4 88.8 71.0 76.50
Llama3(8B), SPruFT, $r=128$ 24.49GB 159.4M(1.98%) 87.6 77.4 71.4 85.4 70.2 79.8 90.8 81.8 80.55
Llama3(8B), LoRA, $r=16$ 24.88GB 32.5M(0.41%) 83.8 83.6 73.8 86.6 67.4 79.6 91.2 78.6 80.58
Llama3(8B), SPruFT, $r=32$ 21.37GB 27.3M(0.34%) 83.2 82.6 66.6 86.0 65.6 82.4 89.8 80.4 79.58
Llama3(8B), LoRA, $r=32$ 25.28GB 65.0M(0.81%) 83.0 68.8 71.4 85.2 66.2 81.0 91.6 79.6 78.35
Llama3(8B), SPruFT, $r=64$ 21.81GB 54.5M(0.68%) 84.4 78.0 71.0 83.2 69.0 79.0 84.4 83.4 79.05
Llama3(8B), LoRA, $r=64$ 26.04GB 130.0M(1.62%) 86.2 81.6 67.8 81.8 66.0 73.8 86.2 78.2 77.73
Llama3(8B), SPruFT, $r=128$ 22.71GB 109.1M(1.36%) 89.4 79.0 71.2 86.2 71.8 83.0 86.8 76.2 80.45
Table 15: Similar to Table 6, this is the full table with different rank settings. We also present the results of freezing queue-, key-, and value-projection in this table ().
Model, ft setting BoolQ PIQA SIQA HS WG ARC-c ARC-e OBQA Avg